# On the Extension of Inner Derivations from Dense Ideals in Banach Algebras

### 10 Mar 2026 v1

@@ Abstract.
Let $A$ be a Banach algebra and $I$ a dense ideal in $A$. A natural question in the theory of operator algebras is whether the property that all derivations $D:A\to I$ are inner (implemented by elements in $I$) implies that all derivations $D:A\to A$ are inner (implemented by elements in $A$). We present a rigorous negative answer to this question. By utilizing the algebra of compact operators $A=K(H)$ and the dense ideal of finite-rank operators $I=F(H)$ on a separable infinite-dimensional Hilbert space $H$, we demonstrate that while every derivation into $F(H)$ is inner, there exist outer derivations on $K(H)$. Furthermore, we generalize this result to Schatten $p$-classes and discuss the cohomological implications and the role of approximate identities.

### 14 Mar 2026 v2

@@ Abstract.
Let $A$ be a Banach algebra and $I$ a dense ideal in $A$. A natural question in the theory of operator algebras is whether the property that all derivations $D:A\to I$ are inner (implemented by elements in $I$) implies that all derivations $D:A\to A$ are inner (implemented by elements in $A$). We present a rigorous negative answer to this question. By utilizing the algebra of compact operators $A=K(H)$ and the dense ideal of finite-rank operators $I=F(H)$ on a separable infinite-dimensional Hilbert space $H$, we demonstrate that while every derivation into $F(H)$ is inner, there exist outer derivations on $K(H)$. Furthermore, we generalize this result to Schatten $p$-classes and discuss the cohomological implications and the role of approximate identities. Moreover, the main results and counterexamples presented in this paper have been formally verified using the Lean theorem prover.
@@ 7. Cohomological Interpretation
8. Formal Verification
To ensure the strict logical rigor of the topological and algebraic obstructions detailed in this work, the primary results have been formally verified using the Lean 4 theorem prover. The machine-checked proofs rely on the foundational functional analysis and operator algebra frameworks provided by mathlib [8].
The formalization confirms the spatial restrictions on derivations mapping into $F(H)$ and $\mathcal{S}_{p}(H)$, as well as the explicit construction of the outer derivations on $K(H)$ via the Calkin algebra and commutator bounds. The complete Lean source code, including all definitions, lemmas, and exact theorem statements corresponding to this paper, is publicly available in the accompanying repository at: https://github.com/tavakkoliamirmohammad/inner-deriv-dense-lean.git.